OpenStudy (anonymous):

The supplement of an angle is 9 times greater than the measure of the complement of the angle. Find the measure of the angle.

OpenStudy (anonymous):

I think the answer could be 81.

OpenStudy (anonymous):

c=complement s=supplement x=angle

OpenStudy (anonymous):

angle + supplement = 180 => angle =180 - supplement angle + complement = 90 => angle = 90 -compplement 180 - supp = 90 - comp ==> => supp -comp = 90 supp = 9comp 9comp - comp = 90 => 8comp = 90 comp = 90 /8 =11.25 angle = 78,75

OpenStudy (anonymous):

c=90-x 9c=180-x s=9c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the system must reolver

OpenStudy (anonymous):

I subtract the two equations to obtain the complementary angle

OpenStudy (anonymous):

so did u ever get the answer